OpenStudy (anonymous):

Simplify. y 5 ÷ y 3 · y 2

mathslover (mathslover):

\[\huge{y^2*y^2=y^4}\]

OpenStudy (anonymous):

\\[y ^{(5-3+2)}=y ^{4}\]

OpenStudy (anonymous):

SO WE DONT FOLLOW ORDER OF OPERATIONS?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I mean PEMDAS, because then multiplication is before division :S

OpenStudy (anonymous):

÷; ;*are the same priority

OpenStudy (anonymous):

???

OpenStudy (anonymous):

you don't need care which one is first.

OpenStudy (anonymous):

so can u explain me this why n^6 · n^5 · n^4 ÷ n^3 · n^2 ÷ n = n^13

OpenStudy (anonymous):

n(6+5+4-3-2-1) = n^9 :S

OpenStudy (anonymous):

!!!!!!+2

OpenStudy (anonymous):

i feel stupid haha thanks

OpenStudy (anonymous):

thanks for ur patience :)

OpenStudy (anonymous):

never mind it